OpenStudy (mackenzie2013):

Laura is designing a rectangular patio in her garden. She has enough tiles to cover 270 square feet. She wants the length of the patio to be 17 feet longer than the width. What dimensions should Laura use for the patio? l = 26 ft; w = 9 ft l = 27 ft; w = 10 ft l = 22 ft; w = 5 ft l = 35 ft; w = 18 ft

OpenStudy (jiteshmeghwal9):

Let the required width be w then length=w+17 Area=Length*width=w(w+17)=270

OpenStudy (mikey):

I would multiply each answer for example 26*9 and see which one equals 270ft but make sure the width is 17feet shorter than the length

563blackghost (563blackghost):

:P beat me to it